# Particle Trajectory Representation Learning with Masked Point Modeling

### v2

@@ 4 Method
Figure 2 outlines PoLAr-MAE’s pretraining, adapted from Point-MAE with an added energy reconstruction task. The LArTPC image is converted into a point cloud, processed through a modified grouping module. Some groups are masked, while visible ones are embedded into latent tokens. A ViT-based autoencoder, with a heavy encoder and light decoder, reconstructs full patches. For energy reconstruction, a permutation-equivariant embedding module processes 3D point positions, and a linear head predicts energy using decoded masked embeddings and encoded 3D positions.
@@ 4.1.1 Patch grouping
The common method for grouping point clouds into smaller patches involves iterative farthest point sampling (FPS) to sample group centers, followed by a $k$-nearest neighbors ($k$-NN) or ball query to find nearby neighbors for inclusion in the patch. However, in our dataset, where point density varies in regions containing signal, traditional FPS combined with $k$-NN or ball queries often proves insufficient. Specifically, these approaches result in either an excessive number of ungrouped points or excessive overlap between local patches (see Figure 3).
@@
Volumetric point cloud grouping.
To address this issue, we introduce a volumetric sampling method, centrality-based non-maximum suppression (C-NMS), which extends the conventional NMS approach to operate on spherical regions. C-NMS enables efficient sampling of the minimal number of patches while maintaining low levels of ungrouped points and controlled overlap between patches. The method proceeds as follows:
@@
    A large set of candidate group centers is sampled, e.g., via FPS or another method.
@@
The resulting groups form a variable number $G$ of point groups, each containing a variable number $K$ of points. This approach approximates the effect of voxelizing the volume spanned by all points with a voxel pitch equal to the sphere diameter, while inherently reducing aliasing artifacts. Unlike FPS+$k$-NN, which lacks control over patch overlap, C-NMS ensures a tunable overlap between patches. We argue that $k$-NN grouping post-C-NMS is worse than a ball query, as the latter naturally constrains the domain of points within a canonical radius, effectively mitigating outlier inclusion within patches. Additional commentary on this approach, including the pseudocode implementation, can be found in Appendix B.
Dynamic patch grouping.
In traditional FPS+{$k$-NN, ball query}, researchers often employ a fixed number of patches and a fixed number of points per patch (e.g., someone pre-training on ShapeNet (Chang et al., 2015) may produce 128 patches with 32 points each). By contrast, C-NMS dynamically determines both the number of tokens $G$ and the number of points per token $K$ based on two parameters: the sphere radius $r$ and the overlap factor $f$. The computational overhead of the additional NMS step involves an extra ball query on seed points, which can be efficiently executed using CUDA-accelerated libraries such as pytorch3d (Ravi et al., 2020), but is certainly non-negligible.
Patch normalization.
After grouping via C-NMS, each patch is normalized by subtracting the patch center: $\mathbf{x}_{i}\rightarrow\mathbf{x}_{i}-\mathbf{c},$ where $\mathbf{x}_{i}$ is the original point position’s first three dimensions, and $\mathbf{c}$ is the patch center. Following PointNeXT (Qian et al., 2022), we scale each normalized point by the sphere radius $r$, such that all patch points’ positional coordinates lie within $[-1,1]^{3}$: $\mathbf{x}_{i}\rightarrow\frac{\mathbf{x}_{i}}{r}.$ Note that we only apply these normalizations to the coordinates of each point, and not the energy value.
Mini-PointNet for patch embedding.
After grouping, each patch group is encoded into a single latent vector using a permutation-invariant mini-PointNet (Qi et al., 2017a).
Transformer encoder.
The patch embeddings from the mini-PointNet form the input sequence to a vanilla Transformer encoder, which computes global contextual information across all patch tokens. To preserve spatial information lost during grouping, positional embeddings are added to the patch tokens. Each positional embedding is learned and corresponds to the 3D coordinates of the patch center. We omit the final LayerNorm (Ba et al., 2016) operation in the Transformer encoder, as it resulted in better representations.
@@ 4.2 Patch-based masked event modeling
To facilitate effective self-supervised pretraining for point cloud data, we employ a masked autoencoder framework. A masking module selects a random subset of patch tokens, leaving the remaining unmasked tokens to be processed through the patch embedding module and Transformer encoder, producing contextually enriched embeddings. A shallow decoder then takes the encoded visible tokens along with a duplicated number of a placeholder learnable mask token. Positional embeddings corresponding to the spatial position of the patch centers are added to both visible and masked token embeddings, enabling the decoder to reconstruct the latent representations of the masked tokens.
4.2.1 Full Patch Reconstruction
@@
The Point-MAE approach takes the decoded mask tokens and applies a simple linear layer to predict the original masked point groups. Formally, for each masked token, the decoder outputs $\textbf{x}_{g}\in\mathbb{R}^{K_{\text{max}}\times(3+1)}$ predicted points, where $K_{\text{max}}$ is the maximum possible number of points in any group. To handle variable group sizes, we take only the first $K$ points for each masked group, where $K$ corresponds to the actual number of points in the group. The reconstruction is evaluated using the Chamfer Distance (CD) loss between the predicted points $\{\mathbf{x}_{g}^{\text{pred}}\}$ and the original points $\{\mathbf{x}_{g}^{\text{true}}\}$ of the masked patches.
4.2.2 Per-point Energy Reconstruction
To further facilitate trajectory representation learning, we introduce an auxiliary energy reconstruction task, whereby the network learns to predict per-point energies given their individual 3D positions.
@@ 4.2.2 Per-point Energy Reconstruction
Standard permutation-invariant PointNets struggle with per-point predictions in variable-sized groups due to unordered processing. We address this with an Equivariant Mini-PointNet that imposes structured order via sinusoidal positional encodings (Vaswani, 2017) processed through a learnable MLP (see right of Fig. 2). Unlike the original, order embeddings are injected after the first shared MLP layer, intentionally breaking permutation invariance while retaining padding invariance (when padding values match group points). A final linear projection layer concatenates positional masked tokens with the predicted mask tokens from the decoder, predicting $K_{\text{max}}$ energy values per group via the $L_{2}$ loss. This enables precise per-point energy reconstruction while handling variable group sizes.
@@ 5 Results
5.1 Pre-training
We train our model on the training set described in §3) using a batch size of 128 for 500,000 iterations, corresponding to about 60 epochs. We use a grouping radius of 5 voxels (15 mm) with an optimized C-NMS overlap fraction $f=0.73$, mask 60% of input tokens randomly, and adopt ViT-S-like specifications for the encoder/decoder. With our dataset, this results in about 400 groups per event (Figure 4). Full architectural details are provided in Appendix C, as well as additional commentary on optimizing our grouping strategy in Appendix B.
@@ 5.1 Pre-training
Figure 5: Visualization of learned representations after pretraining. We use PCA to project the learned representations into RGB space. The Random Initialized model displays a strong positional bias, while the PointMAE model shows bias for individual trajectories. The PointMAE model contains visually similar embeddings as PoLAr-MAE, and is not pictured. Best viewed zoomed in. Additional examples (with embeddings for PointMAE) can be found in Appendix D.
Table 1 provides SVM F-scores (F1) – defined as the harmonic mean of the precision and recall of the classifier, i.e. $F_{1}=\frac{2}{\text{recall}^{-1}+\text{precision}^{-1}}$ – at the end of training for both the PointMAE and PoLAr-MAE. We give metrics for PointMAE from two training runs – one with 100k training steps ($\sim$2 epochs), and one with 500k training steps ($\sim$60 epochs). These metrics make clear that the model definitively learns the semantic meaning of different types of trajectories without explicitly providing particle ID information. By just pre-training, the SVM can be taught to classify tokens as containing tracks and showers with F-scores of 99.4% and 97.7%, respectively. Particle types that often span one to two tokens are much less discernible, and as such are harder, but not impossible, to discern, having F-scores of 51.8% for Michels and 44.0% for delta rays . Interestingly, the model learns to very effectively discriminate between tracks and showers just within around 10 epochs of training.
Table 1: SVM Validation Results. We present final SVM validation F1 scores for each class—where F1 is the harmonic mean of precision and recall—and an overall mean across classes. The best results are highlighted.
Per-class F1
Method Steps F1 Track Shower Delta Michel
PointMAE 100k 0.680 0.990 0.970 0.436 0.325
PointMAE 500k 0.719 0.993 0.967 0.572 0.342
PoLAr-MAE 500k 0.732 0.994 0.977 0.518 0.440
To get a qualitative understanding of what these representations of trajectories work out, we plot out the center of each token in an event and color it according to a casting of its latent vectors to RGB space using PCA. Shown in Figure 5, the pretrained models clearly shows an understanding of semantics between tracks, showers, deltas, and Michels. A comparison to the latent representations of a model initialized with random weights is shown for comparison. Indeed, even trajectories who cross very close to one another, despite being very spatially close, have different representations. Tokens along a single particle trajectory are shown to have similar representations. Additional examples can be found in the appendix.
Hyperparameters.
The critical hyperparameters we use include a group radius of 5 voxels (15 mm) with C-NMS overlap fraction $f=0.73$, oversampling size of 2048, and a context length of 512 tokens. Groups are capped at 32 points, with farthest point sampling (FPS) applied when exceeding this limit. We mask 60% of input tokens randomly during training and adopt ViT-S-like specifications for encoder/decoder. Chamfer distance serves as the training objective. Full architectural details and secondary hyperparameters are provided in Appendix C.
5.2 Semantic Segmentation
Event 1 Event 2 Event 3
Truth
Refer to caption Refer to caption Refer to caption
PoLAr-MAE PEFT
Refer to caption Refer to caption Refer to caption
PoLAr-MAE FFT
Refer to caption Refer to caption Refer to caption
Figure 6: Visualization of finetuned PoLAr-MAE models on the semantic segmentation task. Classes are one of {track , shower , delta ray , Michel . Strong performance for the PoLAr-MAE models trained with parameter-efficient fine-tuning (PEFT) methods and via full fine-tuning (FFT) are shown. Most of the error is from incorrectly classifying spatially small semantic types, e.g., delta rays alongside tracks. The finetuned PointMAE predictions are visually similar to that of the PoLAr-MAE models, and are not pictured. Additional examples (with predictions from the PointMAE models) can be found in Appendix E. Best viewed zoomed in.
To evaluate the effectiveness of the representations learned by the PoLAr-MAE encoder, we fine-tune the model to perform dense classification of individual voxels as one of the four particle types: tracks , shower , delta rays , and Michel . We freeze the encoder and train simple feature upsampling and classification heads to predict per-point classes, following (Qi et al., 2017b). We fine-tune both PoLAr-MAE and PointMAE pretrained models on 10,000 events, 10x less than the number of events used to train SPINE (Drielsma et al., 2021), and use the focal loss (Lin et al., 2018) along with additional loss weights proportional to the inverse frequencies of each class to balance the loss across classes. We perform full fine-tuning (optimizing all parameters of the model) as well as linear probing, i.e., freezing all components of the model except the feature upscaling and segmentation heads, to evaluate how well the representations the encoder learns actually are.
Table 2: Semantic segmentation results. We present F1 scores for semantic segmentation of each class within the validation set – where F1 is the harmonic mean of precision and recall—and an overall mean across classes. The best results are highlighted.
Per-class F1
Model F1 Track Shower Delta Michel
PointMAE PEFT 0.772 0.965 0.983 0.569 0.572
PoLAr-MAE PEFT 0.798 0.961 0.990 0.542 0.698
PointMAE FFT 0.831 0.963 0.994 0.561 0.807
PoLAr-MAE FFT 0.837 0.964 0.994 0.569 0.823
Table 2 shows results on semantic segmentation across the two pretraining strategies and the two fine-tuning strategies using F1 scores as evaluation metrics. Overall, our fine-tuned models performs exceptionally well, especially given these models were fine-tuned with just 10,000 events. PoLAr-MAE overall outperforms or is comparable to PointMAE in the FFT mdoels, however there is a clear trade-off in Michel/delta classification with the PEFT models. Qualitative examples of semantic segmentation predictions are shown in Fig. 6. A direct comparison of average precision values reported in (Drielsma et al., 2021) are shown in Figure 7. Critically, our model outperforms the fully supervised approach for shower and track classification by 2% and 0.1%, respectively, while performing fairly worse for Michel and delta-ray classification. A more detailed comparison between models can be found in Appendix E.
Figure 7: Comparison to the fully supervised approach. We compare to reported average classification precision values found in (Drielsma et al., 2021). Notably, our fully fine-tuned PoLAr-MAE model outperforms the Sparse UResNet-based SPINE while training on 10x less labeled data.
6 Discussion
This work demonstrates the viability of SSL for TPC data analysis, bridging the gap between foundational methodologies in computer vision and the unique challenges of high-energy physics. By reducing dependency on labeled data and enabling scalable cross-experiment analyses, our approach represents a small step toward the development of generalized, domain-agnostic foundation models for particle physics.
An important aspect of this work is the introduction of the Centrality-Based Non-Maximum Suppression (C-NMS) tokenization approach. Unlike traditional grouping methods such as farthest point sampling (FPS) with $k$-nearest neighbors (k-NN) or ball query, C-NMS dynamically adjusts to the spatial density of TPC point clouds. This ensures minimal overlap and maximized coverage of particle trajectories, addressing key limitations of existing methods.

### v3

@@ 4 Method
Figure 2 outlines PoLAr-MAE’s overall architecture and pretraining strategy, adapted from Point-MAE with an added energy reconstruction task. The LArTPC image is converted into a point cloud and processed through a modified grouping module. Some groups are masked, while visible ones are embedded into latent discrete tokens. A Vision Transformer (ViT) (Ranftl et al., 2021)-based autoencoder, with a heavy (i.e., many parameters) encoder and light (i.e., comparatively less parameters) decoder, reconstructs full patches. For energy reconstruction, a permutation-equivariant embedding module processes 3D point positions, and a linear head predicts per-point energy using decoded masked embeddings and encoded 3D positions.
@@ 4.1.1 Patch Grouping
The common method for grouping point clouds into smaller patches involves farthest point sampling (FPS) – an iterative algorithm that selects points that are farthest away from the previously selected points – to sample group centers, followed by a $k$-nearest neighbors ($k$-NN) or ball query to find nearby neighbors for inclusion in the patch. However, in our dataset, where point density varies in regions containing signal, traditional FPS combined with $k$-NN or ball queries proves insufficient. Specifically, these approaches result in either an excessive number of ungrouped points or excessive overlap between local patches (see Figure 3).
@@
To address this issue, we introduce a volumetric sampling method, centrality-based non-maximum suppression (C-NMS), which extends the conventional NMS approach used in computer vision to operate on spherical regions. C-NMS enables efficient sampling of the minimal number of patches while maintaining low levels of ungrouped points and controlled overlap between patches. The method proceeds as follows:
@@
    A large set of candidate group centers is sampled, e.g., via FPS or by taking the entire point cloud.
@@
The resulting groups form a variable number $G$ of point groups, each containing a variable number $K$ of points. This approach approximates the effect of voxelizing the volume spanned by all points with a voxel pitch equal to the sphere diameter, while inherently reducing aliasing artifacts. Unlike FPS+$k$-NN, which lacks control over patch overlap, C-NMS ensures a tunable overlap between patches. In traditional FPS+{$k$-NN, ball query}, researchers often employ a fixed number of patches and a fixed number of points per patch. By contrast, C-NMS dynamically determines both the number of tokens $G$ and the number of points per token $K$ based on two parameters: the sphere radius $r$ and the overlap factor $f$. This method, which is highly parallelizable, has been implemented as a custom CUDA kernel for extremely fast processing. Additional commentary, further information, and ablation studies on this approach can be found in Appendix B.
After grouping via C-NMS, each patch is normalized by subtracting the patch center: $\mathbf{x}_{i}\rightarrow\mathbf{x}_{i}-\mathbf{c},$ where $\mathbf{x}_{i}$ is the original point position’s first three dimensions, and $\mathbf{c}$ is the patch center, defined by the sphere centroid returned by C-NMS. Following PointNeXT (Qian et al., 2022), we scale each normalized point by the sphere radius $r$, such that all patch points’ positional coordinates lie within $[-1,1]^{3}$: $\mathbf{x}_{i}\rightarrow\frac{\mathbf{x}_{i}}{r}.$ Note that we only apply these normalizations to the coordinates of each point, and not the energy value.
4.1.2 Patch Encoding
After grouping, each patch group is encoded into a single latent vector using a permutation-invariant mini-PointNet (Qi et al., 2017a). The mini-PointNet is a small MLP network that first uplifts each individual point in each group to a higher latent dimension and then performs max-pooling along the channel dimension, so as to “collapse” all individual point embeddings into a single latent vector. Since groups may contain less than $K$ points per patch, we pad each group with copies of its first point to avoid propagating padding artifacts. The mini-PointNet is invariant to input permutations, and as such is also invariant to this form of padding.
The patch embeddings from the mini-PointNet form the input sequence to a vanilla Transformer encoder (Vaswani, 2017), which computes global contextual information across all patch tokens. To preserve spatial information lost during grouping, positional embeddings are added to the patch tokens. Each positional embedding (Vaswani, 2017) is learned via a three-layer MLP and corresponds to a mapping from the 3D coordinates of the patch center to some latent vector. We omit the final LayerNorm (Ba et al., 2016) operation in the Transformer encoder, as it resulted in better representations.
@@ 4.2 Patch-based masked event modeling
To facilitate effective self-supervised pretraining for LArTPC data, we employ a masked autoencoder framework. A masking module selects a random subset of patch tokens, leaving the remaining unmasked tokens to be processed through the patch embedding module and Transformer encoder, producing contextually enriched embeddings. A shallow decoder then takes the encoded visible tokens along with a duplicated number of placeholder learnable mask tokens. Positional embeddings corresponding to the spatial position of the patch centers are added to both visible and masked token embeddings, enabling the decoder to reconstruct the latent representations of the masked tokens given spatial cues. This process is shown in Fig. 2.
@@
The Point-MAE framework takes the decoded mask tokens and applies a simple linear layer to predict the original masked point groups. Formally, for each masked token, the decoder outputs $\textbf{x}_{g}\in\mathbb{R}^{K_{\text{max}}\times(3+1)}$ predicted points, where $K_{\text{max}}$ is the maximum possible number of points in any group. To handle variable group sizes, we take only the first $K$ points for each masked group, where $K$ corresponds to the number of points in the ground truth group. The reconstruction is evaluated using the Chamfer Distance ($d_{CD}$) loss between the predicted points $\{\mathbf{x}_{g}^{\text{pred}}\}$ and the original points $\{\mathbf{x}_{g}^{\text{true}}\}$ of the masked patches, which is defined as
$d_{CD}(S_{1},S_{2})=\sum_{x\in S_{1}}\min_{y\in S_{2}}||x-y||_{2}^{2}+\sum_{y\in S_{2}}\min_{x\in S_{1}}||x-y||^{2}_{2}.$ (2)
This loss is the sum of the minimum distance between each point in one set $S_{1}$ with any point in another $S_{2}$ plus its permutation, and is permutation invariant to point ordering.
@@
To further facilitate trajectory representation learning, we introduce an auxiliary energy reconstruction task, whereby the network learns to predict per-point energies given their individual 3D positions. This is a natural extension, as the energy deposition over the length of a given trajectory $(\frac{dE}{dx})$ is an important discriminator for particle identification (PID) in any LArTPC analysis. However, permutation-invariant PointNets struggle with per-point predictions in variable-sized groups due to unordered processing. We address this with an Equivariant Mini-PointNet that imposes structured order via sinusoidal positional encodings (Vaswani, 2017) processed through a learnable MLP. Unlike the original, positional embeddings corresponding to the (arbitrary) tensor ordering of points within each patch are injected after the first shared MLP layer, intentionally breaking permutation invariance while retaining the aforementioned padding invariance. A final linear projection layer concatenates positional masked tokens with the predicted mask tokens from the decoder, predicting $K_{\text{max}}$ energy values per group via the $L_{2}$ loss. This enables precise per-point energy reconstruction while handling variable group sizes.
4.3 Pre-training
We train our model on the training set described in Sec. 3 using an effective batch size of 512 across 4 NVIDIA A100 GPUs for 500,000 iterations, corresponding to about 60 epochs. We use a grouping radius of 5 voxels (15 mm) with an optimized C-NMS overlap fraction $f=0.73$, mask 60% of input tokens randomly, and adopt ViT-S-like specifications (Ranftl et al., 2021) for the encoder/decoder. Groups are capped at 32 points, with farthest point sampling (FPS) (Qi et al., 2017b) applied when exceeding this limit. With our dataset, this results in about 400 groups per event (Figure 4). Full architectural details and secondary hyperparameters are provided in Appendix C.
@@ 5 Results
We evaluate the PoLAr-MAE framework using quantitative metrics to assess the quality of its pre-trained representations and its performance on downstream semantic segmentation. Additionally, we provide qualitative visualizations of the learned features and model attention maps.
5.1 Quantitative Results
5.1.1 Linear Token Classification
Table 1 provides SVM F-scores (F1), defined as the harmonic mean of the precision and recall of the classifier (i.e., $F_{1}=\frac{2}{\text{recall}^{-1}+\text{precision}^{-1}})$ at the end of training for both the Point-MAE and PoLAr-MAE. We give metrics for Point-MAE from two training runs: one with 100,000 training steps ($\sim$2 epochs), and one with 500,000 training steps ($\sim$60 epochs). These metrics make clear that the model definitively learns the semantic meaning of different types of trajectories without explicitly providing particle ID information. By just pre-training, the SVM can be taught to classify tokens as containing tracks and showers with F-scores of 99.4% and 97.7%, respectively. Particle types that often span one to two tokens are much less discernible, and as such are harder, but not impossible, to discern, having F-scores of 51.8% for Michels and 44.0% for delta rays . Interestingly, the model’s representations of patches contain tracks and showers become easily linearly separable just within around 10 epochs of training.
Table 1: SVM Validation Results. We present final SVM validation F1 scores for each class—where F1 is the harmonic mean of precision and recall—and an overall mean across classes. Bold entries highlight the best results.
Per-class F1
Method Steps F1 Track Shower Delta Michel
Point-MAE 100k 0.680 0.990 0.970 0.436 0.325
Point-MAE 500k 0.719 0.993 0.967 0.572 0.342
PoLAr-MAE 500k 0.732 0.994 0.977 0.518 0.440
5.1.2 Semantic Segmentation
Figure 5: Data efficiency of the pre-trained model. We fine-tune PoLAr-MAE on small labeled datasets consisting of 100, 1,000, and 10,000 events and compare precision results to the supervised UResNet baseline when trained on the same number of events. Despite being fine-tuned on just 100 events, PoLAr-MAE reaches $>0.99\%$ precision for track/shower disambiguation, and outperforms the supervised approach in all three dataset sizes.
Event 1 Event 2 Event 3
Truth
Refer to caption Refer to caption Refer to caption
PEFT
Refer to caption Refer to caption Refer to caption
FFT
Refer to caption Refer to caption Refer to caption
Figure 6: Visualization of finetuned PoLAr-MAE models on the semantic segmentation task. Classes are one of {track , shower , delta ray , Michel }. Strong performance for the PoLAr-MAE models trained with parameter-efficient fine-tuning (PEFT) methods and via full fine-tuning (FFT) are shown. Most of the error results from incorrectly classifying spatially small semantic types, e.g., delta rays alongside tracks. The finetuned Point-MAE predictions are visually similar to that of the PoLAr-MAE models, and are not pictured. Additional examples (with predictions from the Point-MAE models) can be found in Appendix E. Best viewed zoomed in.
To evaluate the effectiveness of the representations learned by the PoLAr-MAE encoder, we fine-tune the model to perform dense classification of individual voxels as one of the four particle types: tracks, showers, delta rays, and Michel. We train simple feature upsampling and classification heads to predict per-point classes, following PointNet++ (Qi et al., 2017b) and Point-MAE (Pang et al., 2022). We fine-tune both PoLAr-MAE and Point-MAE pretrained models on 10,000, 1,000, and 100 events, which are 10x, 100x, and 1000x less than the number of events used to train the supervised baseline UResNet (Drielsma et al., 2021b). We opt to use the Focal loss (Lin et al., 2018), which is defined as
$\text{FL}(p_{t})=-\alpha_{t}(1-p_{t})^{\gamma}\log{p_{t}},$ (3)
where $p_{t}\in[0,1]$ is the model’s estimated probability for the ground truth class, $\alpha_{t}$ is the inverse frequency of the true label within the entire dataset, and $\gamma=2$. The Focal loss, an extension of standard cross entropy loss, focuses the relative loss towards harder (low confidence) examples by downweighting the loss if the model is very confident in its prediction. We apply weights $\alpha_{t}$ to compensate for the large class imbalance between tracks/showers and Michels/deltas.
We perform full fine-tuning (optimizing all parameters of the model) as well as parameter-efficient fine-tuning (PEFT), i.e., freezing all components of the model except the relatively tiny feature upscaling and segmentation heads, to evaluate how well the representations the encoder learns actually are.
5.1.3 Data efficient fine-tuning
To further evaluate the effectiveness of the representations learned by the PoLAr-MAE encoder, we evaluate the model when fine-tuned for semantic segmentation on an extremely small number of labeled events – 100 and 1000, and 10,000. To demonstrate the power of initializing the segmentation model with pre-trained weights, we compare segmentation results to SPINE’s UResNet baseline trained from scratch on the same amounts of data. As shown in Figure 5, PoLAr-MAE achieves a precision of 0.995 for showers and 0.993 for tracks after being fine-tuned on just 100 LArTPC events. This compares favorably to the fully supervised UResNet, which when trained on 100 events yields a precision of 0.326 for showers and 0.933 for tracks. Despite comparatively lower precision scores on delta ray and Michel electron classification overall, PoLAr-MAE generally outperforms or matches the fully supervised UResNet across these categories at all three smaller data sizes, notably for Michels with 10,000 events.
Table 2: Semantic segmentation results. We present F1 scores for semantic segmentation of each class within the validation set – where F1 is the harmonic mean of precision and recall – and an overall mean across classes. The best results are highlighted.
Per-class F1
Model F1 Track Shower Delta Michel
Point-MAE PEFT 0.772 0.965 0.983 0.569 0.572
PoLAr-MAE PEFT 0.798 0.961 0.990 0.542 0.698
Point-MAE FFT 0.831 0.963 0.994 0.561 0.807
PoLAr-MAE FFT 0.837 0.964 0.994 0.569 0.823
Table 2 shows results on semantic segmentation across the two pretraining strategies and the two fine-tuning strategies using F1 scores as evaluation metrics. Overall, our fine-tuned models perform exceptionally well, especially given these models were fine-tuned with just 10,000 events. PoLAr-MAE overall outperforms or is comparable to Point-MAE in the FFT models, however there is a clear trade-off in Michel/delta classification with the PEFT models. Qualitative examples of semantic segmentation predictions are shown in Fig. 6. A direct comparison of precision values reported in (Drielsma et al., 2021b) are shown in Figure 7. Critically, our model outperforms the fully supervised approach for shower and track classification by 2% and 0.1%, respectively, while performing fairly worse for Michel and delta-ray classification. A more detailed comparison between models can be found in Appendix E.
Figure 7: Comparison to the fully supervised approach. We compare PoLAr-MAE fine-tuned on 10,000 events to reported average classification precision values of SPINE trained for 100,000 events (Drielsma et al., 2021b). Notably, our fully fine-tuned PoLAr-MAE model outperforms the Sparse UResNet-based SPINE on track and shower classification while training on 10x less labeled data.
5.2 Qualitative Results
5.2.1 PCA of Patch Features
To get a qualitative understanding of what these representations of trajectories work out, we plot out the center of each token in an event and color it according to a casting of its latent vectors to RGB space using principal component analysis (PCA). To ensure that the colors more directly relate to the underlying semantic meaning of individual tokens, we regress out any spatial bias by training a simple affine model to predict embeddings from their patch centers. A full explanation of this process is described in Appendix D.1.
Shown in Figure 8, the pretrained models clearly shows an understanding of semantics between tracks, showers, deltas, and Michels. A comparison to the latent representations of a model initialized with random weights (also de-biased) is shown for comparison. Indeed, even trajectories that cross very close to one another, despite being very spatially close, have different representations. Tokens along a single particle trajectory are shown to have similar representations. Particle trajectories that all come from the same vertex also differ in representations. Additional examples can be found in the appendix.
@@ 5.2.1 PCA of Patch Features
Figure 8: Visualization of spatially debiased learned representations after pretraining. We use PCA to project the learned representations into RGB space after spatial debiasing. The randomly initialized model displays a strong positional bias, while the Point-MAE model shows a clearer bias towards individual trajectories. The Point-MAE model contains visually similar embeddings as PoLAr-MAE, and is not pictured. Best viewed zoomed in. Additional examples (with embeddings for Point-MAE) as well as the debiasiing procedure can be found in Appendix D.
Figure 9: Visualization of attention maps from the 6th head of the 5th encoder layer. Top row: full event used as input (left) and zoomed-in regions showing attention focusing on small trajectories (right). Bottom four panels: attention patterns when querying tokens (red dots) on longer tracks and a shower. Attention largely conforms to individual particle instances. Color intensity reflects the strength of attention contributed by other tokens to the queried token.
5.2.2 Attention Maps
To gain qualitative insights into the representations learned by PoLAr-MAE, particularly how the model processes spatial relationships between different parts of an event, we visualize the attention mechanisms within the Transformer encoder. Figure 9 showcases these attention patterns from the 6th attention head in the 5th layer of the encoder. In these visualizations, a specific token (the “query point,” marked in red) is selected, and all other tokens in the point cloud are colored according to the attention strength they contribute to this queried token. High attention (yellow) indicates tokens that the model deems most relevant to understanding or contextualizing the queried token.
The figure is structured to demonstrate this phenomenon across different scales and particle types. The top row first presents the full input event alongside several zoomed-in regions. These magnified views highlight how the attention mechanism can delineate even small, distinct trajectories, such as short tracks or Michel electrons (e.g., top right panels, within colored bounding boxes). When querying tokens on longer tracks or within electromagnetic showers (bottom four panels of Fig. 9, the attention patterns consistently concentrate along the queried particle’s trajectory or within the characteristic spread of the shower.
This consistent focusing of attention along individual particle trajectories, whether they are short and isolated or long and part of a complex event, shows that the self-supervised pre-training task encourages the model to learn not just local geometric features but also a more high-level understanding of particle contiguity and separation. This intrinsic grouping capability is a highly desirable property for downstream tasks, as it suggests the learned representations are inherently structured to distinguish between different particle instances.
6 Future Work and Discussion
This work presents the first successful implementation and adaptation of a self-supervised, masked autoencoding framework, PoLAr-MAE, for representation learning directly on the sparse, 3-D point cloud data from Liquid Argon Time Projection Chambers (LArTPCs). Our results demonstrate that by learning to reconstruct masked portions of particle trajectories, the model develops a robust and transferable understanding of the underlying physics without relying on any labeled examples. These learned representations are highly effective, proving amenable to abstract downstream tasks and enabling remarkable data efficiency.
Successfully transferring SSL paradigms from computer vision to the HEP domain often necessitates domain-specific adaptations. Our development of the Centrality-Based Non-Maximum Suppression (C-NMS) tokenization serves as an example of a modification. Unlike traditional grouping methods common in computer vision, C-NMS is designed to dynamically adjust to the characteristic spatial density variations within TPC point clouds, ensuring minimal token overlap and maximized coverage of particle trajectories. Such adaptations are crucial when applying existing architectures to new scientific domains with distinct data characteristics.